A chemist mixes distilled water with a 85 percent solution of sulfuric acid to produce a 45 percent solution. If 7 liters of distilled water are used, how much 45 percent solution is produced?
Answer by TimothyLamb(4379)

---
x = liters of 85% solution
y = liters of distilled water
---
85x + 0y = 45(x + y)
y = 7
---
put the system of linear equations into standard form:
---
85x = 45x + 45y
y = 7
---
40x - 45y = 0
y = 7
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = liters of 85% solution = 7.875
y = liters of distilled water = 7
---
system matrix
40 -45
0 1
---
inverse of system matrix
0.025 1.125
0 1
---
determinant of system matrix = 40
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php